An Olympic diver has a mass of 45.8 kg and has 7129 J of gravitational potential energy.
What is the height of the diving board? Place your answer in the box below to 3 significant digits.
1
Expert's answer
2020-05-25T11:18:23-0400

The potential energy in the gravitational field is given by:


"W = mgh"

where "m = 45.8 kg" is the mass of the body and "h" is the height above the Earth surface. Let's express the "h" from this formula and substitute numerical values.:


"h = \dfrac{W}{mg} = \dfrac{7129J}{45.8kg\cdot 9.81 m/s^2} = 15.867m"

Answer. h = 15.867 m.